OpenStudy (anonymous):

What value is a discontinuity of x squared plus 8 x plus 4, all over x squared minus x minus 6?

hero (hero):

You are given \[\frac{x^2 + 8x + 4}{x^2 - x - 6}\] and asked to find the value of the discontinuity, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-1 -3 -3 or -2

OpenStudy (anonymous):

how do you find the discontinuity

hero (hero):

The discontinuity of the rational expression occurs where the numerator and denominator has common factors, correct?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

So its -2

hero (hero):

And if we factor the rational expression completely we should be able to find those common expressions right?

OpenStudy (anonymous):

Hmm, yes.

hero (hero):

How did you arrive at -2? Would you mind explaining?

hero (hero):

Because \(x^2 + 8x + 4\) doesn't appear to be factorable. Did you make a mistake while posting the expression?

OpenStudy (anonymous):

Because it goes into 8 and 4?

hero (hero):

-2 is correct, however, that's not the reason why.

hero (hero):

\(x^2 + 8x + 4\) factors, just not over integers. Observe that \(\dfrac{x^2 + 8x + 4}{x^2 - x - 6} \) factors to \(\dfrac{(x - 2(\sqrt{3}-2))(x + 2(\sqrt{3} +2))}{(x - 3)(x + 2)}\)